Suppose a four period weighted average is being used to forecast demand. Weights for the periods are: wt-4 = 0.05 wt-3 = 0.15 wt-2 = 0.35 wt-1 = 0.45 Demand observed in the previous four periods was: At-4 = 10245 At-3 = 9800 At-2 = 9796 At-1 = 9714 What will be the demand forecast for period t? (Keep one decimal place in your answer).
Answer #1

Solution :-

Demand forecast for period t

= (Wt-4 * At-4) + (Wt-3 * At-3) + (Wt-2 * At-2) + (Wt-1 * At-1)

= (0.05*10245) + (0.15*9800) + (0.35*9796) + (0.45*9714)

= 512.25 + 1470 + 3428.6 + 4371.3

= 9782.15

DEMAND FORECADT FOR PERIOD T = 9782.2